DIRC Operations Weekly Report

Period: Saturday 2006/07/28 - Friday 2006/08/04

Oncall: Rolf Andreassen
Fast Monitoring: Bradley Wogsland, Rolf Andreassen

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060804.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060804 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 10.8*1033 (previous week: 10.6*1033)

BaBar integrated lumi: 5.24 fb-1 (previous week: 3.11 fb-1)

Run 5b PEP delivery: 85.54 fb-1

Run 5 total delivery: 144.38 fb-1 Run 5 total BaBar recorded: 139.35 fb-1 including 11.62 fb-1 off-peak data

PEP activities this week:

  • Off-peak running
  • Delivery reaches 400/fb! (By PEP statistics. BaBar disagrees - but maybe next week.)
  • Fast dithering hardware up, software still being worked on.

    PEP activities planned next week :

  • Run off-peak until the 8th, then begin MD preparation.

    Foreseen work opportunities (PEP down periods):

  • PEP will do several things in preparation for down; schedule here.

    BaBar activities/issues this week:

    DIRC activities

    See DIRC logbook for details

  • Sat 29th : There was a failure in the power to the degasing pump panel. Since the SOB water level monitor is on the same panel, this caused SOB dump and SOB water level alarms. Jerry was able to restore the power and fix the problem; he also moved the degasing pump over to UPS power again, to relieve the load on the normal power.
  • Thu 3rd : Background monitoring plots have stopped updating - hardcoded 'End of Run 5'? Await Nicolas' return.
  • Thu 3rd : There is a bit of a funny smell in the DIRC gas rack. Jerry will investigate.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Manpower situation

    Links